OpenStudy (anonymous):

WILL FAN AND MEDAL What is the slope of the line 4x – 8y – 16 = 0 in simplest form?

OpenStudy (anonymous):

@fallenangelorchid pls help

OpenStudy (kamibug):

4x - 8y - 16 = 0 You can change this to slope-intercept form... Add 16 to both sides. 4x - 8y = 16 Subtract 4x from both sides. -8y = -4x + 16 Divide both sides by -8. y = 1/2x + 4 Now it's in the form y = mx + b where m is the slope! :)

OpenStudy (fallenangelorchid):

She deserves a medal :D

OpenStudy (anonymous):

she does and she got one

OpenStudy (anonymous):

so the slope \[\frac{ 1 }{ 2 }\]

OpenStudy (kamibug):

\(\huge\cal\color{royalblue}{Correct!~:)}\)

OpenStudy (anonymous):

then thats the awnser

OpenStudy (kamibug):

You got it. ^.^

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

hey can you help me with one more

OpenStudy (kamibug):

Sure, and I messed up, sorry, Lol, the last term is -2 xD Not 4. :P

OpenStudy (anonymous):

\[y-1=\frac{ 4y-2 }{ 3 }\]

OpenStudy (anonymous):

we have to solve it any ideas

OpenStudy (kamibug):

Multiply both sides by 3. 3y - 3 = 4y - 2 Add 3 to both sides. 3y = 4y + 1 Subtract 4y from both sides. -y = 1 Multiply both sides by -1. y = -1

OpenStudy (anonymous):

thank you so much you are amazing

OpenStudy (kamibug):

My pleasure! :)